A bus traveled 387 km in 5h. One hour of the trip was in city traffic. The bus's city speed was just half of its speed on open highway. The rest of the trip was on open highway. Find the bus's city speed.
Answer by sudhanshu_kmr(1152)

let bus's city speed = x
speed on open highway = 2x
total distance traveled = x*1 + 2x * 4 = 9x
according to question
9x = 387
=> x = 387/9 = 43


bus's city speed = 43 km/h